# SmartWilds: Multimodal Wildlife Monitoring Dataset

### v1

@@ Abstract
We present the first release of SmartWilds, a multimodal wildlife monitoring dataset. SmartWilds is a synchronized collection of drone imagery, camera trap photographs and videos, and bioacoustic recordings collected during summer 2025 at The Wilds safari park in Ohio. This dataset supports multimodal AI research for comprehensive environmental monitoring, addressing critical needs in endangered species research, conservation ecology, and habitat management. Our pilot deployment captured four days of synchronized monitoring across three modalities in a 220-acre pasture containing Pere David’s deer, Sichuan takin, Przewalski’s horses, as well as species native to Ohio, including bald eagles, white-tailed deer, and coyotes. We provide a comparative analysis of sensor modality performance, demonstrating complementary strengths for landuse patterns, species detection, behavioral analysis, and habitat monitoring. This work establishes reproducible protocols for multimodal wildlife monitoring while contributing open datasets to advance conservation computer vision research. Future releases will include synchronized GPS tracking data from tagged individuals, citizen science data, and expanded temporal coverage across multiple seasons.
@@ 1 Introduction
Conservation biology requires comprehensive ecosystem monitoring to inform evidence-based management decisions, yet traditional approaches provide fragmented views of wildlife activity and habitat use. The integration of multiple sensing modalities, powered by edge AI and computer vision approaches offers transformative opportunities for automated environmental monitoring at unprecedented scales (Besson et al.,, 2022; Tuia et al.,, 2022; Pringle et al.,, 2025; Kline et al.,, 2025). Our research advances multimodal AI for wildlife monitoring by creating datasets and models that enable object detection and tracking across environmental conditions, support fine-grained species classification from multi-sensor data, and facilitate behavioral analysis from long video sequences.
We present a multimodal dataset from sensor deployments at The Wilds Conservation Center (The Wilds,, 2025)  [^1] https://imageomics.github.io/naturelab/. Our dataset was curated to evaluate multi-sensor fusion techniques and benchmark machine learning approaches for conservation applications across visual, acoustic, and environmental data streams. This dataset poses challenges to existing computer vision methods through its combination of rare species detection, variable conditions, and the need for behavioral state recognition over extended time periods. This work supports research areas including endangered species monitoring, conservation and restoration ecology, and habitat assessment for both exotic endangered species and native biodiversity conservation. Our contributions include: (1) a synchronized multimodal ecological dataset with comprehensive metadata, (2) reproducible protocols for environmental monitoring sensor networks, and (3) support for sensor fusion to advance multimodal learning research in environmental contexts.
@@
Figure 1: Representative images and data study design. GPS and time-stamp metadata allow for cross-referencing between modalities. (a) Diagram of dataset modalities, citizen science images, GPS tags, acoutic data, camera trap and drone images, joined via location and time-stamp metadata. (b) Example of multi-modal data cross-referencing using metadata. Camera trap view (TW02) of Pere David’s deer synchronized with drone image of Pere David’s deer wading in lake.
@@ 2 Related Work
Recent advances have introduced multimodal wildlife monitoring datasets addressing diverse conservation challenges (Smith and Pinter-Wollman,, 2021; Buxton et al.,, 2018; Besson et al.,, 2022). MammAlps (Gabeff et al.,, 2025) combines multi-view video with synchronized audio data to analyze wild mammal behaviors in alpine environments, demonstrating the power of multimodal approaches for detailed behavioral analysis. BuckTales (Naik et al.,, 2024) provides multi-UAV tracking data for wild antelope identification and re-identification, advancing techniques for individual animal monitoring. The PanAf-FGBG dataset (Brookes et al.,, 2025) explores how environmental backgrounds impact wildlife behavior recognition, while YOLO-Behaviour (Chan et al.,, 2025) demonstrates automated behavior quantification frameworks. KABR (Kholiavchenko et al.,, 2024) focuses on Kenyan wildlife behavior recognition from drone footage, contributing to conservation AI in African ecosystems. However, most existing datasets focus on specific taxonomic groups or behavioral tasks rather than comprehensive ecosystem monitoring. Our work contributes synchronized multi-sensor data collection protocols designed specifically for conservation digital twin development, with planned integration of GPS tracking data for individual-level behavioral analysis.
@@ 3.1 Study Site
The Wilds is a 10,000-acre conservation center in southeastern Ohio, home to endangered species conservation programs and native wildlife restoration efforts (The Wilds,, 2025). Our pilot deployment focused on a single 220-acre enclosure containing a breeding population of GPS-tagged Pere David’s deer (Elaphurus davidianus), chosen specifically to enable future integration with individual tracking data. Sensor sites were selected chronologically based on observed wildlife activity patterns and strategic coverage of diverse habitat types within the study pasture, illustrated in Fig. 2 and described in Table 3. Camera trap sites prioritized high deer activity areas, particularly around water sources, while bioacoustic monitors targeted diverse acoustic environments from open grasslands to woodland edges. Where possible, existing structures were utilized for sensor mounting to minimize environmental impact and maximize equipment protection.
@@ 5.2 Comparison of Sensor Modalities
We summarize the relative performance each sensing modality across eight key performance dimensions relevant to conservation monitoring applications (Table 2). These metrics were selected based on established frameworks for conservation technology evaluation and practical deployment considerations in wildlife monitoring contexts. Spatial range and resolution determines monitoring coverage and detection capabilities across different habitat scales (Tuia et al.,, 2022; Pringle et al.,, 2025). Temporal range and resolution captures both short-term behavioral events and long-term ecological patterns Besson et al., (2022). Species detectability captures different sensor modalities ability to sense specific taxonomic groups, especially more cryptic species such as birds or insects,(Smith and Pinter-Wollman,, 2021). Behavioral detail is important for understanding complex social interactions of group-living animals and individual to group-level responses to environmental changes (Kline et al.,, 2025). Deployment effort and data volume captures practical considerations affecting scalability and cost of long-term monitoring efforts (Besson et al.,, 2022).
Table 2: Comparative analysis of sensor modality performance across key conservation monitoring metrics. Metrics selected based on established frameworks for wildlife monitoring technology evaluation (Tuia et al.,, 2022; Besson et al.,, 2022) and practical deployment considerations in conservation settings. Performance rating: Poor, Moderate, Good.
@@ 7 Data Availability Statement
Dataset cards with representative samples and comprehensive metadata are currently available for review: https://zenodo.org/records/17183089. The complete dataset will be made publicly available on Hugging Face upon publication of this manuscript under a CC0-1.0 license.
@@
We thank Dan Beetem, The Wilds Conservation Center and the Columbus Zoo for their support in facilitating this project. All data collection, including drone flights, was conducted under supervision of The Wilds Animal Welfare team, with permission from The Wilds Scientific Advising Board.
This project is supported by U.S. National Science Foundation under Award No. 2112606, Intelligent Cyberinfrastructure with Computational Learning in the Environment (ICICLE), and Award No. 2118240, HDR Institute: Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning.

### v2

@@ Abstract
We present the first release of SmartWilds, a multimodal wildlife monitoring dataset. SmartWilds is a synchronized collection of drone imagery, camera trap photographs and videos, and bioacoustic recordings collected during summer 2025 at The Wilds safari park in Ohio. This dataset supports multimodal AI research for comprehensive environmental monitoring, addressing critical needs in endangered species research, conservation ecology, and habitat management. Our pilot deployment captured four days of synchronized monitoring across three modalities in a 220-acre pasture containing Pere David’s deer, Sichuan takin, Przewalski’s horses, as well as species native to Ohio. We provide a comparative analysis of sensor modality performance, demonstrating complementary strengths for landuse patterns, species detection, behavioral analysis, and habitat monitoring. This work establishes reproducible protocols for multimodal wildlife monitoring while contributing open datasets to advance conservation computer vision research. Future releases will include synchronized GPS tracking data from tagged individuals, citizen science data, and expanded temporal coverage across multiple seasons.
@@ 1 Introduction
Conservation biology requires comprehensive ecosystem monitoring to inform evidence-based management decisions, yet traditional approaches provide fragmented views of wildlife activity and habitat use. The integration of multiple sensing modalities, powered by edge AI and computer vision approaches offers transformative opportunities for automated environmental monitoring at unprecedented scales (bessontoward2022; tuiaperspectives2022; pringleopportunities2025; kline2025studying). Our research advances multimodal AI for wildlife monitoring by creating datasets and models that enable object detection and tracking across environmental conditions, support fine-grained species classification from multi-sensor data, and facilitate behavioral analysis from long video sequences.
We present a multimodal dataset from sensor deployments at The Wilds Conservation Center (thewilds_website)  [^1] https://imageomics.github.io/naturelab/. Our dataset was curated to evaluate multi-sensor fusion techniques and benchmark machine learning approaches for conservation applications across visual, acoustic, and environmental data streams. This dataset poses challenges to existing computer vision methods through its combination of rare species detection, variable conditions, and the need for behavioral state recognition over extended time periods. This work supports research areas including endangered species monitoring, conservation and restoration ecology, and habitat assessment for both exotic endangered species and native biodiversity conservation. Our contributions include: (1) a synchronized multimodal ecological dataset with comprehensive metadata, (2) reproducible protocols for environmental monitoring sensor networks, and (3) support for sensor fusion to advance multimodal learning research in environmental contexts.
@@
Figure 1: Representative images and data study design. GPS and time-stamp metadata allow for cross-referencing between modalities. (a) Diagram of dataset modalities, citizen science images, GPS tags, acoustic data, camera trap and drone images, joined via location and time-stamp metadata. (b) Example of multi-modal data cross-referencing using metadata. Camera trap view (TW02) of the Pere David’s deer synchronized with the drone image of the Pere David’s deer wading in the lake.
@@ 2 Related Work
Recent advances have introduced multimodal wildlife monitoring datasets addressing diverse conservation challenges (smithobserving2021; buxtonpairiningcamera2018; bessontoward2022). MammAlps (gabeff2025mammalps) combines multi-view video with synchronized audio data to analyze wild mammal behaviors in alpine environments, demonstrating the power of multimodal approaches for detailed behavioral analysis. BuckTales (naik2024bucktales) provides multi-UAV tracking data for wild antelope identification and re-identification, advancing techniques for individual animal monitoring. The PanAf-FGBG dataset (brookes2025panaf) explores how environmental backgrounds impact wildlife behavior recognition, while YOLO-Behavior (chan2025yolo) demonstrates automated behavior quantification frameworks. KABR (kholiavchenko2024kabr) focuses on Kenyan wildlife behavior recognition from drone footage, contributing to conservation AI in African ecosystems. However, most existing datasets focus on specific taxonomic groups or behavioral tasks rather than comprehensive ecosystem monitoring. Our work contributes synchronized multi-sensor data collection protocols designed specifically for conservation digital twin development, with planned integration of GPS tracking data for individual-level behavioral analysis.
@@ 3.1 Study Site
The Wilds is a 10,000-acre conservation center in southeastern Ohio, home to endangered species conservation programs and native wildlife restoration efforts (thewilds_website). Our pilot deployment focused on a single 220-acre enclosure containing a breeding population of GPS-tagged Pere David’s deer (Elaphurus davidianus), chosen specifically to enable future integration with individual tracking data. Sensor sites were selected chronologically based on observed wildlife activity patterns and strategic coverage of diverse habitat types within the study pasture, illustrated in Fig. 2 and described in Table 3. Camera trap sites prioritized high deer activity areas, particularly around water sources, while bioacoustic monitors targeted diverse acoustic environments from open grasslands to woodland edges. Where possible, existing structures were utilized for sensor mounting to minimize environmental impact and maximize equipment protection.
@@ 5.2 Comparison of Sensor Modalities
We summarize the relative performance each sensing modality across eight key performance dimensions relevant to conservation monitoring applications (Table 2). These metrics were selected based on established frameworks for conservation technology evaluation and practical deployment considerations in wildlife monitoring contexts. Spatial range and resolution determines monitoring coverage and detection capabilities across different habitat scales (tuiaperspectives2022; pringleopportunities2025). Temporal range and resolution captures both short-term behavioral events and long-term ecological patterns bessontoward2022. Species detectability captures different sensor modalities ability to sense specific taxonomic groups, especially more cryptic species such as birds or insects,(smithobserving2021). Behavioral detail is important for understanding complex social interactions of group-living animals and individual to group-level responses to environmental changes (kline2025studying). Deployment effort and data volume captures practical considerations affecting scalability and cost of long-term monitoring efforts (bessontoward2022).
Table 2: Comparative analysis of sensor modality performance across key conservation monitoring metrics. Metrics selected based on established frameworks for wildlife monitoring technology evaluation (tuiaperspectives2022; bessontoward2022) and practical deployment considerations in conservation settings. Performance rating: Poor, Moderate, Good.
@@ 7 Data Availability Statement
The dataset is freely available on HuggingFace under a CC0-1.0 license. The dataset is arranged by modality, including camera trap images, bioacoustic recordings, and drone video footage.
Future releases will include GPS tracking data of Pere David’s deer, citizen science images captured using mobile phones and cameras, and detailed weather and habitat data from satellite data sources. We also plan to expand citizen science integration through crowdsourced imagery validation and expert-guided annotation workflows.
@@
We thank Dan Beetem, The Wilds, and the Columbus Zoo & Aquarium for their support in facilitating this project. All data collection, including drone flights, was conducted under the supervision of the Director of Animal Management, with permission from The Wilds Animal Care and Use Committee.
This project is supported by the AI Institute for Intelligent Cyberinfrastructure with Computational Learning in the Environment (ICICLE), the Imageomics Institute, and the AI and Biodiversity Change (ABC) Global Center. ICICLE is funded by the US National Science Foundation under Award No. 2112606, the Imageomics Institute is funded by the US National Science Foundation’s Harnessing the Data Revolution (HDR) program under Award No. 2118240 (Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning). The ABC Global Center is funded by the US National Science Foundation under Award No. 2330423 and the Natural Sciences and Engineering Research Council of Canada under Award No. 585136. This work draws on research supported by the Social Sciences and Humanities Research Council.